Monochromatic light from a distant source is incident on a slit 0.750 mm wide. On a screen 2.00 m away, the distance from the central maximum of the diffraction pattern to the first minimum is measured to be 1.35 mm. Calculate the wavelength of the light.

1
Expert's answer
2022-01-12T08:31:43-0500

Solution:

Conditions for the position of the minima is,

"\tan \theta=\frac{y}{D}"

Rearrange the above equations to get angle of diffraction,

"\theta=\tan ^{-1}\left(\frac{y}{D}\right)"

Substitute "1.35 \mathrm{~mm} for\ \mathrm{y} and\ 2.00 \mathrm{~m}\ for\ D."

"\begin{aligned}

\theta &=\tan ^{-1}\left(\frac{(1.35 \mathrm{~mm})\left(\frac{10^{-3} \mathrm{~m}}{1 \mathrm{~mm}}\right)}{2.00 \mathrm{~m}}\right) \\

&=0.038674^{\circ} \\

& \approx 0.0387^{\circ}

\end{aligned}"

Expression for the wavelength of the light is,

"\lambda=\frac{d \sin \theta}{n}"

Substitute d for "0.750 \mathrm{~mm}, 0.0387^{\circ} for\ \theta, and\ 1 for\ n" .

"\begin{aligned}

\lambda &=\frac{(0.750 \mathrm{~mm})\left(\frac{10^{-3} \mathrm{~m}}{1 \mathrm{~mm}}\right) \sin \left(0.0387^{\circ}\right)}{1} \\

&=506.58 \times 10^{-9} \mathrm{~m} \\

& \approx 507 \mathrm{~nm}

\end{aligned}"

The wavelength of the light is "507 \mathrm{~nm}."